# Higher-Order Energy-Decreasing Exponential Time Differencing Runge–Kutta methods for Gradient Flows

### v1

@@ 3 Examples of energy decreasing ETDRK schemes
$\phi_{0}(z)=e^{z},\quad\phi_{k+1}=\frac{\phi_{k}(z)-\phi_{k}(0)}{z},\quad\phi_{k}(0)=1/k!$ (3.1)
@@ 3.2 Second-order ETDRK scheme
The first term is positive-definite since $(z+1/\phi_{1})$ is positive and $E_{L}$ is positive-definite, the second term is also positive-definite because $\phi_{1}>\phi_{2}>0$, and the third term is obviously positive-definite. Therefore, $\Delta$ is positive-definite and the ETDRK2 scheme is unconditionally energy stable. ∎
@@ 3 $\sum_{i=1}^{s}b_{i}J\psi_{2,i}=0$
and $J$ denotes arbitrary bounded operators. In particular, the classical ETDRK3 scheme below from Cox and Mathhews [5] does not satisfy the conditions of our theorem (see Fig. 2) so it is not unconditionally energy decreasing.
@@
To begin with, we first present numerical evidences. We plot in Fig. 2 the smallest eigenvalues of $\frac{1}{2}(\Delta+\Delta^{T})$ for the three schemes above, and we observe that the smallest eigenvalues of (3.10) and (3.11) are positive, indicating that the conditions of our theorem are satisfied, while the smallest eigenvalues for the scheme (3.9) when $Z$ approaches zero become negative, which indicates that the conditions of our theorem are violated. To rigorously show that the above two ETDRK schemes decrease the energy, we only need to prove the positive-definiteness of $\Delta^{\prime}$ for the above schemes, which could be verified by checking all leading principal minors.
@@
For the schemes (3.10) and (3.11), the determinants of all leading principal minors of Δsuperscriptnormal-$\Delta^{\prime}$ are positive, and thus Δsuperscriptnormal-$\Delta^{\prime}$ is positive definite. Thus, the schemes (3.10) and (3.11) unconditionally decrease the energy for general gradient flows (2.7).
@@
According to (3.10), we derive the expressions for determinants of leading principal minors
@@
Similarly we could define the non-exponential term $P_{3}$ and the exponential term $R_{3}$ for $z^{6}\text{Det}(\Delta)$, and for $z<-6$, we have $P_{3}(z)>P_{3}(-6)=148.5$, $|R_{3}(z)|<86$ and thus $z^{6}\text{Det}(\Delta^{\prime})>P_{3}(z)-|R_{3}(z)|>0$. Combine all above results, we have $\text{Det}(\Delta^{\prime}_{2\times 2})>0,\text{and }\text{Det}(\Delta^{\prime})>0$ for all negative $z$.
@@ 4 $\sum_{i=1}^{s}b_{i}c_{i}K\psi_{2,i}=0$
where $\psi_{i}(z)=\phi_{i}(z)-\sum_{k=1}^{s}b_{k}c_{k}^{j-1}/(j-1)!,\psi_{i,j}=\phi_{i}c_{j}^{i}-\sum_{k=1}^{j-1}a_{jk}c_{k}^{i-1}/(i-1)!$, and $J,K$ denote arbitrary bounded operators (for more details, see [15]). These fourth-order conditions are more complicated than previous third-order conditions so that it is more difficult to construct fourth-order ETDRK schemes. One of the most used ETDRK4 schemes is from Cox and Mathhews [5]:
@@
We plot in Fig. 3 the smallest eigenvalues of $\frac{1}{2}(\Delta+\Delta^{T})$ for the above two schemes, the results clearly indicate that the above two schemes do not satisfy our conditions for energy stability.
@@
Moreover, we checked all existing fourth-order ETDRK schemes that we could find, and also tried to search from a wide family of four-stage and five-stage fourth-order ETDRK methods, but none of them satisfy the requirement of our theorem. Therefore, the existence of energy stable ETDRK4 schemes is still an open problem.
@@ 4 Numerical Experiments
In this section we carry out some numerical experiments to illustrate the convergence and energy decay property of ETDRK schemes for different phase-field models. We first verify the temporal convergence rates with smooth initial data for the Allen–Cahn and Cahn–Hilliard equations. Next we study the large time step behavior for the Allen–Cahn, Cahn–Hilliard and phase field crystal equations. Then we present energy curves with different time steps and stabilization constants to observe their influence. Finally we present an adaptive time stepping example which makes full use of the unconditional energy stability.
In all simulations, we consider models with periodic conditions, and use a Fourier-spectral method in space with a sufficiently fine mesh so that the spatial discretization errors can be ignored compared with temporal discretization errors. The ETDRK3 scheme (3.10) is used if not specified otherwise. Besides, we also consider a truncated double-well potential $\tilde{F}(u)$ so that it naturally satisfies the Lipschitz condition. More precisely, for a sufficiently large $M$ ($M=2$ is enough for cases in this paper), we replace $F(u)=\frac{1}{4}(u^{2}-1)^{2}$ by
@@ 4.2 Large time step tests
In this subsection we study the accuracy of solutions with large time steps. We set $N=128$, $T=8$, $\epsilon=0.1$ for the Allen–Cahn equation and $\epsilon=0.5$ for the Cahn–Hilliard equation and observe the relation between the error and the time steps $\tau=2^{1-k},k=0,1,..,7$ in Fig. 4. We could see that with large time steps the solutions still have good performances.
We also simulated the phase field crystal model with $N=256$, $\epsilon=0.025$ and $\beta=3$ in $\Omega=(0,32)\times(0,32)$. We use both ETDRK2 and ETDRK3 to solve it with the initial condition $u_{0}=\sin(\frac{\pi x}{16})\sin(\frac{\pi x}{16})$, and computed the solution at $T=1$ with $\tau=2^{-k},k=0,1,...11$. The results are plotted in Fig. 5. We observe that both ETDRK schemes performed well with large time steps. More precisely, with $\tau=0.1$, the errors of ETDRK3 and ETDRK2 are about $O(10^{-4})$ and $O(10^{-3})$, respectively.
@@ 4.3 Energy curves with different time steps and stabilization constants
Now we study how energy curves behave with different time steps and stabilization constants. It is well-known that stabilization may affect the dynamics and cause serious time delay phenomena when lower-order schemes are used or large time steps are taken. We first test the Cahn–Hilliard equation with $N=128$, $\epsilon=0.1$ and $\beta=2$. The results are presented in Fig. 6. We observe that when time steps are large, the time delay phenomena is obvious, while when $\tau\leq 0.01$, the influence becomes negligible.
@@
Next we examine the effect of different stabilization constants. Although theoretically they have to be larger than the Lipschitz constant, we could still run the simulation with smaller stabilization constants. While smaller stabilization constants cause less time delay, but solutions may lose its accuracy and even blow-up. Fig.7 shows that when $\tau=0.1$, the solutions are all inaccurate and have obvious time delay, and the solution with $\beta=0.1$ even has unwanted oscillations. This evidence indicates that the stabilization does help the numerical solutions to maintain stability. When $\tau=0.01$, all three curves with $\beta=0.1,1,2$ are very close, which indicates that the stabilization does not affect much when the time step is small.
@@ 4.4 Long time behavior of the phase field crystal model
Next, we simulate the long time evolution of the phase field crystal (PFC) model with $N=256$, $\epsilon=0.025$, $\Omega=(0,128)\times(0,128)$, and the initial condition
@@ 4.5 Adaptive time stepping
Solutions of gradient flows may vary drastically during some short time intervals, but only slightly at other times. A main advantage of unconditional energy stable schemes is that they can be easily used to construct an adaptive time stepping algorithm, in which the time step is only dictated by accuracy rather than by stability. There are often essential difficulties in applying adaptive time-stepping strategies to other schemes, since most of them do not have robust unconditionally stability with variable step sizes. This is also where the significance of the high-order unconditionally energy stable schemes lies.
For gradient flows, there are several adaptive time stepping strategies, see [2, 13, 23, 36]. Here we make use of the strategy in [24] summarized in the following Algorithm 1. In Step 4 and 6, the time step size which needs to update is given by the formula
@@
along with the restriction of the minimum and maximum time steps. In the above formula, $\rho$ is a default safety coefficient, $tol$ is a reference tolerance to be set, $e$ is the relative error computed at each time level in Step 3 and $r$ is the adaptive rate to be set. In our numerical example, we set $\rho=0.9$ and $tol=5*10^{-3}$, the minimum time step to be $10^{-4}$ and the maximum time step to be $\tau=10^{-2}$ for Fig. $6$ and Fig. $7$ and $\tau=10^{-1}$ for Fig. $8$ and $9$. The initial time step is taken as the minimum time step.
@@ 5 Concluding Remarks
We derived in this paper an general framework to construct arbitrary high order and unconditional energy stable ETDRK schemes for a class of gradient flows, and identified a set of conditions that need to be satisfied for a ETDRK scheme to be unconditional energy stable. In particular, we showed that commonly used third-order and fourth-order ETDRK schemes are not unconditional energy stable, and constructed new third-order ETDRK schemes which are unconditional energy stable.
To the best of our knowledge, this is the first result on higher than second-order unconditional energy stable ETDRK schemes. Here are some extensions and future work to consider:
  • Higher-order: While we provided a set of conditions that ETDRK schemes need to satisfy in order to be unconditional energy stable. But due to the complexity of these conditions, the existence of higher than third-order ETDRK unconditionally energy stable schemes is still an open question. One possibility is to add one more stage which will lead to more free parameters to choose from.
  • Other models: We restricted ourselves to a class of gradient flows with Lipschitz nonlinearities. Since the ETDRK schemes are based on the Duhamel’s Principle so the only error comes from the numerical integration, we expect that the new ETDRK schemes behave well for many other models with even more complicated physical structures, although proving rigorously unconditional energy stability for more complicated models would be a challenge.
We are hopeful that our results in this paper can be extended to more general systems and to higher than third-order ETDRK schemes.
Acknowledgements
J. Shen and Z. Fu are partially supported by NSFC 12371409. J. Yang is supported by the National Natural Science Foundation of China (NSFC) Grant No. 12271240, the NSFC/Hong Kong RGC Joint Research Scheme (NSFC/RGC 11961160718), and the Shenzhen Natural Science Fund (No. RCJC20210609103819018).

### v2

@@ 3 Examples of energy decreasing ETDRK schemes
$\phi_{0}(z)=e^{z},\quad\phi_{k+1}=\frac{\phi_{k}(z)-\phi_{k}(0)}{z},\quad\phi_{k}(0)=\frac{1}{k!}.$ (3.1)
@@ 3.2 Second-order ETDRK scheme
The first term is positive-definite since $(z+1/\phi_{1})$ is positive and $E_{L}$ is positive-definite, the second term is also positive-definite because $\phi_{1}>\phi_{2}>0$, and the third term is obviously positive-definite. Therefore, $\Delta$ is positive-definite, and the ETDRK2 scheme is unconditionally energy stable. ∎
@@ 3 $\sum_{i=1}^{s}b_{i}J\psi_{2,i}=0$
and $J$ denotes arbitrary bounded operators. In particular, the classical ETDRK3 scheme below from Cox and Matthews [5] does not satisfy the conditions of our theorem (see Fig. 2) so it is not unconditionally energy decreasing.
@@
To begin with, we first present numerical evidences. We plot in Fig. 2 the smallest eigenvalues of $\frac{1}{2}(\Delta+\Delta^{T})$ for the three schemes above, and we observe that the smallest eigenvalues of (3.10) and (3.11) are positive, indicating that the conditions of our theorem are satisfied. In contrast, the smallest eigenvalues for the scheme (3.9) become negative when $Z$ approaches zero, indicating that the conditions of our theorem are violated. To rigorously show that the above two ETDRK schemes decrease the energy, we only need to prove the positive-definiteness of $\Delta^{\prime}$ for the above schemes, which could be verified by checking all leading principal minors.
@@
For the schemes (3.10) and (3.11), the determinants of all leading principal minors of $\Delta^{\prime}$ are positive, and thus $\Delta^{\prime}$ is positive definite. Thus, the schemes (3.10) and (3.11) unconditionally decrease the energy for general gradient flows (2.7).
@@
According to (3.10), we derive the expressions for the determinants of leading principal minors as follows
@@
Similarly we could define the non-exponential term $P_{3}$ and the exponential term $R_{3}$ for $z^{6}\text{Det}(\Delta)$, and for $z<-6$, we have $P_{3}(z)>P_{3}(-6)=148.5$, $|R_{3}(z)|<86$ and thus $z^{6}\text{Det}(\Delta^{\prime})>P_{3}(z)-|R_{3}(z)|>0$. Combining all the above results, we have $\text{Det}(\Delta^{\prime}_{2\times 2})>0,\text{and }\text{Det}(\Delta^{\prime})>0$ for all negative $z$.
@@ 4 $\sum_{i=1}^{s}b_{i}c_{i}K\psi_{2,i}=0$
where $\psi_{i}(z)=\phi_{i}(z)-\sum_{k=1}^{s}b_{k}c_{k}^{j-1}/(j-1)!,\psi_{i,j}=\phi_{i}c_{j}^{i}-\sum_{k=1}^{j-1}a_{jk}c_{k}^{i-1}/(i-1)!$, and $J,K$ denote arbitrary bounded operators (for more details, see [15]). These fourth-order conditions are more complicated than previous third-order conditions so that it is more difficult to construct fourth-order ETDRK schemes. One of the most used ETDRK4 schemes is from Cox and Matthews [5]:
@@
We plot in Fig. 3 the smallest eigenvalues of $\frac{1}{2}(\Delta+\Delta^{T})$ for the two schemes above, the results clearly indicate that the above two schemes do not satisfy our conditions for energy stability.
@@
Moreover, we checked all existing fourth-order ETDRK schemes, and also searched from a wide family of four-stage and five-stage fourth-order ETDRK methods, but none of them satisfy the requirement of our theorem. Therefore, the existence of energy stable ETDRK4 schemes is still an open problem.
@@ 4 Numerical Experiments
In this section we carry out some numerical experiments to illustrate the convergence and energy decay property of our new ETDRK schemes for different phase-field models. We first verify the temporal convergence rates using smooth initial data for the Allen–Cahn and Cahn–Hilliard equations. Next, we study the behavior with large time steps for the Allen–Cahn, Cahn–Hilliard, and phase field crystal equations. Then, we present energy curves for different time steps and stabilization constants to observe their influence. Finally, we present an adaptive time-stepping strategy that takes advantage of the unconditional energy stability.
In all simulations, we consider models with periodic conditions, and use a Fourier-spectral method in space with a sufficiently fine mesh so that the spatial discretization errors can be ignored compared with temporal discretization errors. The ETDRK3 scheme (3.10) is used if not specified otherwise. Besides, we also consider a truncated double-well potential $\tilde{F}(u)$ so that it naturally satisfies the Lipschitz condition. More precisely, for a sufficiently large $M$ ($M=2$ is enough for cases used in this paper), we replace $F(u)=\frac{1}{4}(u^{2}-1)^{2}$ by
@@ 4.2 Large time step tests
In this subsection, we study the accuracy of solutions with large time steps. We set $N=128$, $T=8$, $\epsilon=0.1$ for the Allen–Cahn equation and $\epsilon=0.5$ for the Cahn–Hilliard equation. We plot in Fig. 4 the relation between the error and the time steps $\tau=2^{1-k}$ for $k=0,1,...,7$, and observe that the solutions with large time steps still maintain good accuracy.
Next, we simulate the phase field crystal model with $N=256$, $\epsilon=0.025$, and $\beta=3$ in $\Omega=(0,32)\times(0,32)$. Using both ETDRK2 and ETDRK3 with the initial condition $u_{0}=\sin(\frac{\pi x}{16})\sin(\frac{\pi x}{16})$, we compute the solution at $T=1$ with $\tau=2^{-k}$ for $k=0,1,...,11$. The results are plotted in Fig. 5. We observe that both ETDRK schemes work well with large time steps. More precisely, with $\tau=0.1$, the errors of ETDRK3 and ETDRK2 are about $O(10^{-4})$ and $O(10^{-3})$, respectively.
@@ 4.3 Energy curves with different time steps and stabilization constants
Now we study how energy curves behave with different time steps and stabilization constants. It is well-known that stabilization may affect the dynamics and cause serious time delay phenomena when lower-order schemes are used or large time steps are taken. We first test the Cahn–Hilliard equation with $N=128$, $\epsilon=0.1$ and $\beta=2$. The results are presented in Fig. 6. We observe that when time steps are large, the time delay phenomena is obvious, but with smaller time steps $\tau\leq 0.01$, the influence becomes negligible.
@@
Next we examine the effect of different stabilization constants. Although theoretically they have to be larger than the Lipschitz constant, we can still run the simulation with smaller stabilization constants. While smaller stabilization constants cause less time delay, but solutions may lose its accuracy and even blow-up. Fig. 7 shows that when $\tau=0.1$, the solutions are all inaccurate and exhibit obvious time delays, with the solution for $\beta=0.1$ even showing nonphysical oscillations. This evidence indicates that the stabilization does help the numerical solutions to maintain stability. When $\tau=0.01$, all three curves with $\beta=0.1,1,2$ are very close, which indicates that the stabilization has very little effect when the time step is small.
@@ 4.4 Long time behavior of the phase field crystal model
We simulate in this subsection the long time evolution of the phase field crystal (PFC) model with $N=256$, $\epsilon=0.025$, $\Omega=(0,128)\times(0,128)$, and the initial condition
@@ 4.5 Adaptive time stepping
Solutions of gradient flows may vary drastically during some short time intervals, but change only slightly at other times. A main advantage of unconditional energy stable schemes is that they can be easily used with an adaptive time stepping algorithm, in which the time step is only dictated by accuracy rather than by stability. There are often essential difficulties in applying adaptive time-stepping strategies to other schemes, since most of them do not have robust unconditional stability with variable step sizes. This is also where the significance of the high-order unconditional energy stable schemes lies.
For gradient flows, there are several effective adaptive time stepping strategies, see [2, 13, 23, 36]. Here we make use of the strategy in [24] summarized in the following Algorithm 1. In Step 4 and 6, the time step size is given by the formula
@@
along with the restriction of the minimum and maximum time steps. In the above formula, $\rho$ is a default safety coefficient, $tol$ is a reference tolerance, $e$ is the relative error computed at each time level in Step 3, and $r$ is the adaptive rate. In our numerical examples, we set $\rho=0.9$ and $tol=5*10^{-3}$, the minimum time step is $10^{-4}$, while the maximum time step is $\tau=10^{-2}$ for Fig. $6$ and Fig. $7$, and $\tau=10^{-1}$ for Fig. $8$ and $9$. The initial time step is taken as the minimum time step.
@@ 5 Concluding Remarks
In this paper, we presented a general framework for constructing unconditionally energy stable ETDRK schemes of arbitrary order for a class of gradient flows and identified a set of conditions that must be satisfied for an ETDRK scheme to be unconditionally energy stable. In particular, we showed that the widely employed third- and fourth-order ETDRK schemes are not unconditionally energy stable, and constructed new third-order ETDRK schemes which are unconditionally energy stable.
To the best of our knowledge, this is the first rigorous result on unconditionally energy stable ETDRK schemes higher than second-order. Potential future extensions include:
  • Higher than third-order schemes: While we provided a set of conditions that ETDRK schemes need to satisfy in order to be unconditionally energy stable, but due to the complexity of these conditions, the existence of higher than third-order ETDRK unconditionally energy stable schemes is still an open question. One possibility is to add one more stage in the RK part which will lead to more free parameters to choose from.
  • Other models: We restricted ourselves to a class of gradient flows with Lipschitz nonlinearities. Since the ETDRK schemes are based on the Duhamel’s Principle so the only error comes from the numerical integration, we expect that the new ETDRK schemes would behave well for other models with even more complicated physical structures, although proving rigorously unconditionally energy stability for more complicated models would be a challenge.
We are hopeful that our results in this paper can be extended to more general systems and to ETDRK schemes higher than third-order.
Acknowledgments
J. Shen is partially supported by NSFC 12371409. J. Yang is supported by the National Natural Science Foundation of China (NSFC) Grant No. 12271240, the NSFC/Hong Kong RGC Joint Research Scheme (NSFC/RGC 11961160718), and the Shenzhen Natural Science Fund (No. RCJC20210609103819018).